A company has net income of $185,000, a profit margin of 10.00 percent, and an accounts receivable balance of $106,834. Assuming 72 percent of sales are on credit, what is the company's days' sales in receivables?
Answer #1

Total sales = Net income / profit margin

Total sales = 185,000 / 0.1

Total sales = 1,850,000

Credit sales = 72% of 1,850,000 = 1,332,000

Receivables turnover = Net credit sales / receivables

Receivables turnover = 1,332,000 / 106,834

Receivables turnover = 12.4679

Days in sales = 365 / receivables turnover

Days in sales = 365 / 12.4679

Days in sales = 29.28